# Bracketing Transduction Grammar BPE Byte Pair Encoding

### v1

@@ 3.0.4 Facial Expression Generation
3.0.5 Duration Generation
When ground truth timing information is not available we propose a novel sign segmentation approach based on the stitching method described above. Given the ground truth gloss labels, we generate the stitched sequence, $P_{stitch}$, but without step 4 (sign resampling).
Comparing the stitch sequence and the ground truth, we find that the motion can vary due to different lexical variants present compared to our dictionary. However, we find that the handshape is often still consistent. So, we take the keypoints that correspond to the signer’s hands and normalise the rotation so that the index finger metacarpal bone is fixed on the y-axis and the palm is fixed on the xy-plane, giving $P^{H}_{stitch},P^{H}$. Our next step is to align the two sequences so that we can infer the duration of the signs in the ground truth. For this we apply Dynamic Time Warping (DTW), such that;
$A_{i,j}=DTW(P^{H}_{stitch},P^{H})$ (7)
As we know the duration of the isolated signs in the stitched sequence, by analysing the alignment path, $A_{j}$, we can infer the duration of the signs in the original ground truth sequence.
3.0.6 Cutoff Generation
Experiments reveal that each sequence contains a distinct range of frequencies correlated to the signer’s style. Fast motions contain high frequencies, while soft, slow signing involves lower frequencies, typically within the range of 1 to 25 Hz.
@@ 3.0.6 Cutoff Generation
To generate the ground truth cutoffs used in training, we once again apply our stitching approach. For each sequence in the ground truth data, $P$, we produce the equivalent stitched sequence, $P_{stitch}$. We then apply a low-pass filter to $P_{stitch}$ within the range of 1 to 25 Hz and measure the intersection and set difference of the frequencies, denoted as $(FT(P)\cap FT(P_{stitch}))$ and $(FT(P)\setminus FT(P_{stitch}))$, where the Fourier transform is represented as $FT()$. Subsequently, we fit a parametric spline curve to the intersection and set difference. To determine the cutoff we find the frequency that maximises the intersection while minimising the set difference. This provides the cutoff frequency for that sequence. We opt to use this method over just analyzing the frequency in the original sequence as we do not have an ideal filter. Thus, the butterworth filter has unintended effects on the frequency and phase blow the cutoff.
4 Experimental Setup
4.0.1 Implementation Details
In our experiments, we conducted a grid search for optimal hyper-parameters and identified the following settings as the most effective. Our encoder-decoder translation model is constructed with an embedding size of 512 and a feed-forward size of 1024. We find that the optimal number of layers and heads is dataset-dependent, with smaller datasets requiring fewer layers compared to Meine DGS Annotated (mDGS) where we use 3 layers and 4 heads. The models utilise dropout with a probability of 0.1 [41], ReLU activations between layers [1], and pre-layer normalisation for regularisation and training stability. Training employs a ’reduce on plateau’ scheduler with a patience of 5 and a decrease factor of 0.8. The layers are initialised using a Xavier initializer [12] with zero bias, and during training, Adam optimization is employed [19]. The initial learning rate is set to $10^{-4}$, and we train the model until convergence. During decoding, we utilise a greedy search algorithm. The loss scaling factors, $\lambda_{y}$, $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{C}$ are set to 1.0, 0.1, 0.3 and 0.2, respectively. When stitching we enforce a minimum sign length, $U_{min}$ of 4 frames. The cropping threshold, $\alpha_{crop}$ is dataset-dependent, we find values between the range of 0.1 to 0.35 most effective. All sequences are subsampled to 12 frames per second (fps) for computational efficiency.
For each dataset, we create a dictionary of 500 facial expressions. We scale the counter loss by 0.01 and we set an embedding dimension of 512. The encoder and decoder are initialized with the same settings as our translation model.
The angular pose representation comprises 104 angles, while the Euclidean representation consists of 61 keypoints (21 for each hand and 19 for the body and face). The face mesh we add includes 128 keypoints, which are a subset of Mediapipe’s 478-face mesh [26].
For comparison, we train a progressive transformer on each dataset until convergence using the parameters from [38].
4.0.2 Datasets
The approach is tested on three datasets, the Public Corpus of German Sign Language, 3rd release, the mDGS dataset [23], PHOENIX14T [5] and the BSL Corpus T (BSLCPT) [40]. BSLCPT contains 211 participants from 8 regions in the UK, performing 4792 individual signs from a range of age groups. The participants perform narrative, interviews and participate in free conversation. Similarly, mDGS contains 330 participants engaging in free-form signing. Whereas, the PHOENIX14T dataset is extracted from German TV weather broadcasters and contains over 8,000 parallel sequences.
4.0.3 Dictionary
In the following experiment, we test two different dictionaries: 1) collected from isolated examples, and 2) a dictionary created from continuous data. Next, we provide further details about each:
Isolated: Here, the signs are sourced from individuals who perform each sign in isolation, typically starting from and returning to a resting position. When experimenting on the BSLCPT we use the Signbank dataset [7], it contains over 3,000 signs and includes all the lexical variants found in the BSLCPT dataset. However no such dictionary exists for the PHOENIX14T and mDGS dataset, therefore we collect a dictionary from a range of sources such as [23]. We find the mDGS dataset has a target vocabulary of 10,801. However, without the gloss variant, we find the core gloss vocabulary reduces to 4,434. We collect a total of 7,206 signs to experiment with. We use the method described in Sec. 3 (Stitching), step 1 to overcome issues with an incomplete vocabulary. To create word embeddings we apply Facebook’s implementation of Fasttext [27]. When experimenting on the BSLCPT we use the English implementation whereas on PHOENIX14T and mDGS we use the German implementation.
Continuous: Here we create a dictionary using the gloss timing annotations. The signs are taken from the test and dev data only, so that the back translation model has not seen the signs during training. As the examples come from the continuous sign, we omit the cropping step of the stitching pipeline. These dictionaries have an abundance of signs to choose from when stitching. We filter the dictionary and remove short signs as these are most likely co-articulated and therefore not suitable or out of context. We opt to randomly select the first sign in the sequence, and subsequent signs are chosen to ensure the most natural transition. Therefore, we select the sign from the dictionary in which the location of the wrist is closest to the last frame of the previous sign.
4.0.4 Evaluation Metrics
To evaluate our approach we employ a CSLR model (Sign Language Transformers [6]) to conduct back-translation, the same as [14, 36, 39, 52]. A model is trained for each dataset and the parameters are frozen so that results are consistent. BLEU [31], Rouge [25], and chrF [32] scores are computed between the predicted text and the ground truth. To evaluate the pose we employ Dynamic Time Warping Mean Joint Error (DTW-MJE), the metric compresses and stretches time to find the best alignment between the ground truth, $p$, and the predicted pose, $\hat{p}$.
@@ 5.1 Quantitative Evaluation
5.1.1 Text-to-Gloss Translation Results
We start by evaluating the T2G translation performance described in Sec. 3.0.1. Tab. 1 shows the performance on all three datasets. We suggest that the difficulty of a dataset is proportional to the vocabulary and the total number of sequences used in training. We find the best performance on PHOENIX14T data which has the highest number of sequences per token, achieving 18.11 BLEU-4. In comparison to previous works, we find by having the model predict duration, face and cutoff we can achieve higher BLEU-1 scores, but at the cost of a lower BLEU-4 in comparison to [38]. On the more challenging mDGS dataset we find a considerably lower BLUE-4 score due to the larger domain of discourse. The BSLCPT has a smaller domain of discourse in comparison to mDGS but has the fewest sequences per token. Thus, understandably we only achieve a BLEU-4 of 1.67 on the test set. Overall we find the model to be performing as expected.
@@ 5.1.1 Text-to-Gloss Translation Results
PHOENIX14T [38] 55.18 37.10 26.24 19.10 - 54.55 55.65 38.21 27.36 20.23 - 55.41
PHOENIX14T [42] 50.67 32.25 21.54 15.26 - 48.10 50.15 32.47 22.30 16.34 - 48.42
5.1.2 Text-to-Pose Translation Results
Note in the following experiments the back-translation model’s performance (shown as GT top row of Tab. 2, 3 and 4) should be considered the upper limit of performance. In this section, we evaluate the Text-to-Pose (T2P) performance using back translation. To allow for a comparison we train two versions of the PT with the setting presented in [38]. PT is the standard architecture, while PT + GN is trained with Gaussian Noise added to the input. In line with the original paper, we find Gaussian Noise improves the performance, however, our approach still outperforms both models except on DTW-MJE. As discussed previously, other works suffer from regression to the mean caused by the models attempting to minimise their loss function and thus, are incentivised to predict a mean pose. This metric fails to evaluate the content of the sequence, but the higher score does indicate our model is expressive as it is producing sequences further from the mean. For back-translation, we outperform the PT on all metrics. Showing significant improvements in BLEU-1 score of 98% and 269% on the mDGS and BSLCPT dev set (comparing PT + GN and Stitcher (continuous), Tab. 2 and 3).
Deep learning models exhibit a bias toward the data they were trained on and often show poor out-of-domain performance. Unsurprisingly, the performance improves when using the continuous dictionary. We find only a small increase in BLEU-1 of 0.13 on the BSLCPT dev set (Tab. 2), most likely due to the isolated dictionary containing the lexical variants found in the original data. Whereas we see a larger increase on the mDGS dataset (Tab. 3).
@@ 5.1.2 Text-to-Pose Translation Results
PT [38] 0.288 4.40 0.65 0.18 0.00 5.80 8.22 0.292 4.00 0.61 0.10 0.00 5.69 8.02
PT + GN [38] 0.267 4.96 0.55 0.13 0.00 6.38 8.82 0.258 4.47 0.63 0.09 0.00 6.14 8.89
@@
PT [38] 0.229 6.11 0.94 0.21 0.05 8.07 8.36 0.228 6.22 0.98 0.17 0.00 8.23 8.44
PT + GN [38] 0.2245 7.18 1.48 0.40 0.01 8.46 8.38 0.2241 9.22 1.63 0.38 0.01 8.94 8.57
@@
Previous work has primarily focused on G2P translation, therefore to facilitate a meaningful comparison we present two versions of the model. First, a G2P version, where we use the ground truth data and just apply the stitching module, and, secondly our T2P approach (translation then stitching). Results for comparison are provided by [52]. We find our approach outperforms previous work on the BLEU-1 to 2 scores increasing the score by 56% and 19%, respectively (comparing Tab. 4, row 7 and 11). We also find significant improvement in ROUGE and chrF metrics.
Table 4: The results of translating from Gloss-to-Pose (G2P) and Text-to-Pose (T2P) on the RWTH-PHOENIX-Weather-2014T dataset.
@@
PT [38] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
PT + GN [38] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
NAT-AT [14] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
NAT-EA [14] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
PoseVQ-MP [52] 0.146 15.43 10.69 8.26 6.98 - -
PoseVQ-DDM [52] 0.116 16.11 11.37 9.22 7.50 - -
@@
Using a continuous dictionary we can outperform all models except for the VQ based approaches on BLEU-3 to 4. As the VQ model is learning sub-units of a gloss sequence we suggest this gives it an advantage on higher n-grams, as each token can represent multiple signs.
5.1.3 Segmentation Results
To evaluate the stitching segmentation approach we calculate the duration for each gloss in the mDGS dataset test set. We achieve a sign level frame F1-score of 0.6373 a similar score compared to [29] that achieves a top score of 0.63. Demonstrating the effectiveness of stitching for sign segmentation. It is worth noting our approach requires gloss information, but is computationally inexpensive compared to the LSTM used in [29].
@@ 5.2 Qualitative Evaluation
5.2.1 Visual Outputs
To demonstrate the approach’s effectiveness, we present skeleton and video outputs for two sign languages (BSL and DGS) [^1] https://github.com/walsharry/Sign_Stitching_Demos. To enable a fair evaluation, we also share failure cases where incorrect predictions from the initial translation produce repetitive sequences. Furthermore, in the supplementary material, we share visualizations of the produced sequence.
5.2.2 Survey Results
The survey presented a comparison to PT, followed by an ablation of different components of the stitching approach. 17% of people surveyed were native Deaf signers, while 34% were L2 signers or language learners. 87.5% preferred our approach compared to the PT, while the rest selected no preference. 100% of people agreed that applying the filter improved the realism compared to no filtering, while resampling was found to be less important, with 37.5% selecting no preference between the resampled and normal sequence. Interestingly, 41.7% preferred using an isolated dictionary, while 20.8% preferred the continuous dictionary and the rest indicated no preference.

### v2

@@ 3.4 Facial Expression Generation
3.5 Cutoff Generation
@@
To generate the ground truth cutoffs used in training, we once again apply our stitching approach. For each sequence in the ground truth data, $P$, we produce the equivalent stitched sequence, $P_{stitch}$. We then apply a low-pass filter to $P_{stitch}$ within the range of 1 to 25 Hz and measure the intersection and set difference of the frequencies, denoted as $(FT(P)\cap FT(P_{stitch}))$ and $(FT(P)\setminus FT(P_{stitch}))$, where the Fourier transform is represented as $FT()$. Subsequently, we fit a parametric spline curve to the intersection and set difference. To determine the cutoff we find the frequency that maximises the intersection while minimising the set difference. This provides the cutoff frequency for that sequence. We opt to use this method over just analyzing the frequency in the original sequence as we do not have an ideal filter. Thus, the Butterworth filter has unintended effects on frequencies below the cutoff.
@@ 4 Experiments
4.1 Implementation Details
We apply the approach to three datasets, the Public Corpus of German Sign Language, 3rd release, the Meine DGS Annotated (mDGS) dataset [Konrad et al.(2020)Konrad, Hanke, Langer, Blanck, Bleicken, Hofmann, Jeziorski, König, König, Nishio, Regen, Salden, Wagner, Worseck, Böse, Jahn, and Schulder], RWTH-PHOENIX-Weather-2014T (PHOENIX14T) [Camgoz et al.(2018)Camgoz, Hadfield, Koller, Ney, and Bowden] and the BSL Corpus T (BSLCPT) [Schembri(2008)]. To evaluate our approach we employ a CSLR model (Sign Language Transformers [Camgoz et al.(2020)Camgoz, Koller, Hadfield, and Bowden]) to conduct back-translation, the same as [Huang et al.(2021)Huang, Pan, Zhao, and Tian, Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu]. BLEU [Papineni et al.(2002)Papineni, Roukos, Ward, and Zhu], Rouge [Lin(2004)], and chrF [Popović(2015)] scores are computed between the predicted text and the ground truth. Finally, to evaluate the pose we employ Dynamic Time Warping Mean Joint Error (DTW-MJE). In the following experiment, we test two different dictionaries: 1) collected from isolated examples, and 2) a dictionary created from continuous data. Noted as Isolated and continuous in the following tables. Further information about the datasets, dictionaries and model implementation can be found in the supplementary material.
@@ 4.2 Quantitative Evaluation
Text-to-Gloss Translation Results: We start by evaluating the T2G translation performance described in Section 3.1. Table 1 shows the performance on all three datasets. We suggest that the difficulty of a dataset is proportional to the vocabulary and the total number of sequences used in training. We find the best performance on PHOENIX14T data which has the highest number of sequences per token, achieving 18.11 BLEU-4. In comparison to previous works, we find by having the model predict duration, face and cutoff we can achieve higher BLEU-1 scores, but at the cost of a lower BLEU-4 in comparison to [Saunders et al.(2020c)Saunders, Camgoz, and Bowden]. On the more challenging mDGS dataset we find a considerably lower BLUE-4 score due to the larger domain of discourse. The BSLCPT has a smaller domain of discourse in comparison to mDGS but has the fewest sequences per token. Thus, understandably we only achieve a BLEU-4 of 1.67 on the test set. Overall we find the model to be performing as expected.
@@
PHOENIX14T [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 55.18 37.10 26.24 19.10 - 54.55 55.65 38.21 27.36 20.23 - 55.41
PHOENIX14T [Stoll et al.(2018)Stoll, Camgöz, Hadfield, and Bowden] 50.67 32.25 21.54 15.26 - 48.10 50.15 32.47 22.30 16.34 - 48.42
Text-to-Pose Translation Results: Note in the following experiments the back-translation model’s performance (shown as GT top row of Table 2, 3 and 4) should be considered the upper limit of performance. In this section, we evaluate the Text-to-Pose (T2P) performance using back translation. To allow for a comparison we train two versions of the PT with the setting presented in [Saunders et al.(2020c)Saunders, Camgoz, and Bowden]. PT is the standard architecture, while PT + GN is trained with Gaussian Noise added to the input. In line with the original paper, we find Gaussian Noise improves the performance, however, our approach still outperforms both models except on DTW-MJE. As discussed previously, other works suffer from regression to the mean caused by the models attempting to minimise their loss function and thus, are incentivised to predict a mean pose. This metric fails to evaluate the content of the sequence, but the higher score does indicate our model is expressive as it is producing sequences further from the mean. For back-translation, we outperform the PT on all metrics. Showing significant improvements in BLEU-1 score of 98% and 269% on the mDGS and BSLCPT dev set (comparing PT + GN and Stitcher (continuous), Table 2 and 3).
@@
PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.288 4.40 0.65 0.18 0.00 5.80 8.22 0.292 4.00 0.61 0.10 0.00 5.69 8.02
PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.267 4.96 0.55 0.13 0.00 6.38 8.82 0.258 4.47 0.63 0.09 0.00 6.14 8.89
@@
Deep learning models exhibit a bias toward the data they were trained on and often show poor out-of-domain performance. Unsurprisingly, the performance improves when using the continuous dictionary. We find only a small increase in BLEU-1 of 0.13 on the BSLCPT dev set (Table 2), most likely due to the isolated dictionary containing the lexical variants found in the original data. Whereas we see a larger increase on the mDGS dataset (Table 3).
@@
PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.229 6.11 0.94 0.21 0.05 8.07 8.36 0.228 6.22 0.98 0.17 0.00 8.23 8.44
PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.2245 7.18 1.48 0.40 0.01 8.46 8.38 0.2241 9.22 1.63 0.38 0.01 8.94 8.57
@@
PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
NAT-AT [Huang et al.(2021)Huang, Pan, Zhao, and Tian] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
NAT-EA [Huang et al.(2021)Huang, Pan, Zhao, and Tian] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
PoseVQ-MP [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu] 0.146 15.43 10.69 8.26 6.98 - -
PoseVQ-DDM [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu] 0.116 16.11 11.37 9.22 7.50 - -
@@
Table 4: The results of translating from Gloss-to-Pose (G2P) and Text-to-Pose (T2P) on the RWTH-PHOENIX-Weather-2014T test set.
Previous work has primarily focused on G2P translation, therefore to facilitate a meaningful comparison we present two versions of the model. First, a G2P version, where we use the ground truth data and just apply the stitching module, and, secondly our T2P approach (translation then stitching). Results for comparison are provided by [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu]. Note that the previous approaches do not use a dictionary of signs and instead attempt to regress the pose directly from the spoken language. We find our approach outperforms previous work on the BLEU-1 to 2 scores increasing the score by 56% and 19%, respectively (comparing Table 4, row 7 and 11). We also find significant improvement in ROUGE and chrF metrics.
Using a continuous dictionary we can outperform all models except for the Vector Quantisation (VQ) based approaches on BLEU-3 to 4. As the VQ model is learning sub-units of a gloss sequence we suggest this gives it an advantage on higher n-grams, as each token that is predicted can represent multiple signs.
@@ 4.3 Qualitative Evaluation
Visual Outputs: To demonstrate the approach’s effectiveness, we present skeleton and video outputs for two sign languages (BSL and DGS) [^1] https://github.com/walsharry/Sign_Stitching_Demos. Furthermore, in the supplementary material, we share visualizations of the produced sequence.
Survey Results: The survey presented a comparison to PT, followed by an ablation of different components of the stitching approach. 17% of people surveyed were native Deaf signers, while 34% were L2 signers or language learners. 87.5% preferred our approach compared to the PT, while the rest selected no preference. 100% of people agreed that applying the filter improved the realism compared, while resampling was found to be less important, with 37.5% selecting no preference between the resampled and normal sequence.
@@ 5 Conclusion
6 Acknowledgement
We thank Adam Munder, Mariam Rahmani, and Abolfazl Ravanshad from OmniBridge, an Intel Venture. We also thank Thomas Hanke and the University of Hamburg for the use of the Meine DGS Annotated (mDGS) data. This work was supported by Intel, the SNSF project ‘SMILE II’ (CRSII5 193686), the European Union’s Horizon2020 programme (‘EASIER’ grant agreement 101016982) and the Innosuisse IICT Flagship (PFFS-21-47). This work reflects only the author’s view and the Commission is not responsible for any use that may be made of the information it contains.